OpenStudy (anonymous):

write a polynomial function of least degree with integral coefficients that has the given zeros. 2, -1, 3

OpenStudy (ybarrap):

f(x)=(x-2)(x+1)(x-3)= Because f(2)=f(-1)=f(3)=0.

OpenStudy (ybarrap):

Therefore, f(x)=x^3-4 x^2+x+6